OpenStudy (anonymous):

What is the solution set of |2x – 3| – 8 = –1? Answer 5 or –2 No solution 5 or 2 7 or –7

OpenStudy (anonymous):

|2x-3|-8=-1 |2x-3|=7 2x-3=7 or -2x+3=7 2x=10 or -2x=4 x=5 or x=-2

OpenStudy (anonymous):

can you answer this too x < 12 x < –12 or x > 12 x > –12 and x < 12 x > –12

OpenStudy (anonymous):

whats the solution set for it .

OpenStudy (anonymous):

is that the complete question? because x < 12 is just that.

OpenStudy (anonymous):

What is the solution set of |x| > 12?

OpenStudy (anonymous):

|x|>12 x>12 or -x>12 x>12 or x<-12

OpenStudy (anonymous):

it's the first one

OpenStudy (anonymous):

What is the solution set of |2x + 4| < 12? Answer x > –4 and x < 4 x > 8 and x < 4 x < –4 and x > 4 x < 4 and x > –8

OpenStudy (anonymous):

|2x+4|<12 2x+4<12 or -2x-4<12 2x<8 or -2x<16 x<4 or x>-8

OpenStudy (anonymous):

i chose the second one

OpenStudy (anonymous):

ooooohh thanks

OpenStudy (anonymous):

What is the solution set of |4x – 3|+ 2 > 11? Answer x < –1.5 or x > 3 x < –3 or x > 3 x < 1.5 or x > 3 x > –1.5 or x > 3

OpenStudy (anonymous):

i chose the llast one

OpenStudy (anonymous):

|4x-3|+2>11 |4x-3|>9 4x-3>9 or -4x+3>9 4x>12 or -4x>6 x>3 or x<-1.5 you're right

OpenStudy (anonymous):

wait, the inequality changes, so, it's the first one.

OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: y –2x + 3 y < x – 4

OpenStudy (anonymous):

what do you mean with y –2x + 3 there's an inequality missing

OpenStudy (anonymous):

\[\le \]

OpenStudy (anonymous):

you have y ≤ –2x + 3 y < x – 4 so we do y ≤ –2x + 3 < x – 4 and solve –2x + 3 < x – 4 for x –2x + 3 < x – 4 –3x < – 7 x > 2.333 so you can pick any number x greater than 2.333, say 3, plug it into any of the equations y=-2x+3 y=-2(3)+3 y=-6+3 y=-3 so one point that satisfies the inequality is (3,-3).

OpenStudy (anonymous):

but these were my answer choices (–4, 6) (–4, –6) (4, 6) (4, –6)